# Hadwiger Models: Low-Temperature Behavior in a Natural Extension of the Ising Model

### v1

@@ 3 Proving the Phase Diagram
Before we make any specific arguments, first we should note the symmetry of the model. Inverting the state of all hexagons interchanges $E$ vertices for $F$ vertices, $C$ vertices for $H$ vertices, and vice versa. Thus, inverting the hexagons while swapping the energies assigned to the respective vertex states maintains the assignment of probabilities. With this in mind, any statement that concerns particular vertex states applies just as well to its “twin” under inversion symmetry. We will frequently make an argument for one state, then apply the same to its twin by inversion symmetry.
@@
Using techniques from Pirogov–Sinai, we address the large regions dominated by a particular vertex state first, in subsection 3.1. Next, with Slawny’s technique we handle the position of the coexistence lines between these regions in subsection 3.2. Finally, with disagreement percolation and reflection positivity we determine the behavior along the transition lines for which Pirogov–Sinai techniques are inapplicable, in subsection 3.3.
@@ 3.1 Applying Pirogov–Sinai
We begin by describing the ground states of the model, as depicted in Figure 4. Given two vertex states $A,B$, we will refer to the points where $e_{A}=e_{B}$ and both achieve the minimal vertex energy as the $A-B$ transition lines.
Lemma.
Along the $H-F$, $E-C$, or $H-C$ transition lines, there are infinitely many ground configurations. At all other points, there are finitely many ground configurations, all of which are periodic.
@@
First, we consider the case where exactly one vertex state has minimal energy. If a configuration with only that vertex state is possible, it will be minimal. For the $F$ and $E$ vertex states, minimal energy is achieved by the full and empty configurations respectively; any other configuration must have other vertex states, and so the ground configuration is unique and translation invariant. For the $H$ and $C$ vertex states, we start with a single vertex. This could have three distinct assignments, as the $C$ state is realizable by three different three-hexagon configurations. However, given a trio of three hexagons and state $C$, there is only one configuration on its neighbors such that every vertex is in state $C$. By induction, this uniquely determines a periodic ground configuration given one of our original choices, so there are three periodic ground configurations. Along the $E-H$, $E-F$, and $C-F$ transition lines, the minimal vertex states cannot coexist, i.e. neighboring vertices cannot have distinct minimal vertex states, there must be a third state between them. Therefore, all minimal configurations must have only one vertex state, so the set of minimal configurations is the union of the set on the two neighboring regions, and as such is finite and periodic.
Along the $H-F$, $E-C$, and $H-C$ boundary lines, minimal vertex states can coexist. Thus there are infinitely many energy-preserving modifications to particular ground configurations (for example, given the empty configuration, along the $E-C$ line, one can fill in any single hexagon), so the set of ground configurations is infinite. These ground configurations are also not in general periodic: the lower half-plane may have one vertex state, while the upper may have the other.
@@
Lemma.
All Hadwiger models satisfy the Peierls condition, except on the $H-F$, $E-C$, or $H-C$ transition lines.
@@
An $m$-potential is a function on $\omega\in\{0,1\}^{S}$ that can be written as $\sum_{U\subset S}f_{U}(\omega)$, where each $f_{U}$, called a “potential”, depends only on elements in $U$, and there exists a configuration $\omega$ such that $f_{U}$ achieves its minimal value for all $U$. As described above, any Hadwiger Hamiltonian can be represented as the sum over the vertices of terms that depend only on the neighboring hexagons, and may have every vertex in a minimum-energy state, so the Hadwiger models are $m$-potentials. If an $m$-potential’s set of minimal configurations is finite, it satisfies the Peierls condition [FV17]. As just demonstrated, outside the $H-F$, $E-C$, or $H-C$ transition lines the set of minimal configurations is finite, so at those points the Peierls conditions is satisfied
Along the $H-F$, $E-C$, and $H-C$ transition lines, there are infinitely many ground configurations; these are achieved by starting from one ground configuration and flipping any of the infinitely many hexagons that do not increase the energy. Thus, they do not satisfy the Peierls condition and have nonzero entropy at 0 temperature. In particular, the ground states of $H-F$ and $E-C$ are the allowed states of the hard hexagon model, while the ground states of the $H-C$ model are ground states of the frustrated antiferromagnetic Ising model, or equivalently, random lozenge tilings [G21]. ∎
@@
Theorem 1.
At any point such that the unique lowest energy vertex configuration is $H$ or $C$, there exists $T_{0}$ such that for $T<T_{0}$ there are exactly 3 distinct extremal Gibbs states, dominated respectively by the three ground configurations, and every Gibbs state is a linear combination of these.
Proof.
@@
Zahradnik [Z84] states for any periodic Peierls model with finite ground configurations $\exists T_{0}$ such that $\forall T<T_{0}$, every periodic Gibbs state is a linear combination of extremal Gibbs states dominated by a periodic ground configuration. Dobrushin [D85] states that for 2-dimensional models, under the same conditions, every Gibbs state is periodic. Therefore, for any 2-dimensional periodic Peierls model with finite ground configurations $\exists T_{0}$ such that $\forall T<T_{0}$, every Gibbs state is a linear combination of those dominated by the ground states. The model always satisfies the other conditions, so at every Peierls point at sufficiently low temperatures, the space of Gibbs state is exactly those dominated by the ground states.
3.2 Coexistence Lines
@@
At fixed low temperature, there are three coexistence lines to consider, the latter two of which are equivalent under inversion symmetry: $E-F$, $E-H$, and $C-F$. Zahradnik tells us that coexistence lines approach the zero-temperature transition lines, but does not tell us from which direction they approach it, or whether they achieve it at some nonzero temperature. There is a technique by Slawny [S87] to calculate the asymptotic low-temperature positions of these coexistence lines explicitly, by calculating at each point which of the ground configurations has the lowest-energy or highest-multiplicity minimal excitation. We use this technique in a slightly unconventional way. Normally — in two dimensions — Slawny’s technique is used to determine the location of a low-temperature triple point, where three coexistence lines meet, relative to a Peierls zero-temperature triple point. No Peierls triple point exists here, but we can still determine the relative location of the coexistence line by calculating 1-dimensional “slices” perpendicular to each point along the zero-temperature transition line. In each slice, there is a nonzero $T$ such that the transient terms are smaller than the asymptotic term in the expansion we describe below. From Slawny we know the terms vary continuously, so on a closed region there will be a nonzero minimum, such that the overall curve (which by implicit function theorem must be continuous) approximates the asymptotic curve.
Theorem 2.
There exists $T_{0}$ such that $\forall 0<T<T_{0}$, along the $E-H$ transition line, if $e_{F}>e_{C}$, the $E$ state dominates and if $e_{F}<e_{C}$ the $H$ state dominates.
Proof.
@@
Along this line, the $E$ and $H$ vertex states have energy 0, and moving along the line changes the energy of $F$ and $C$. The space of models lies on a sphere, but we’re interested only in the asymptotic shape and position relative to the transition lines, so we can change coordinates at will as long as it does not affect relative position. Project to a plane to simplify coordinates, so along the transition line $e_{F}=f\in[\epsilon,1-\epsilon]$ and $e_{C}=1-f$. Moving perpendicular to this line, $e_{H}=h$. Following Slawny, we can treat the three symmetric ground configurations as a single ground configuration for the purposes of the calculation.
@@
$\dot{P}^{G}(\beta\mathcal{H}_{0}+\mathcal{H})-e_{G}(\mathcal{H}^{\prime})=\dot{P}^{G^{\prime}}(\beta\mathcal{H}_{0}+\mathcal{H}^{\prime})-e_{G^{\prime}}(\mathcal{H}^{\prime})$
where $\dot{P}^{G}$ is the pressure of the gas of excitations from the ground configuration $G$ with respect to a given Hamiltonian, $\mathcal{H}_{0}$ is the Hamiltonian you’re expanding around, $\mathcal{H}^{\prime}$ is the perturbation to the Hamiltonian, and $e_{G}$ is the energy density of $G$ with respect to a given Hamiltonian, equivalent to the energy the Hamiltonian assigns to the associated vertex state. By “pressure of the gas of excitations”, we mean $\log(\sum_{m\in M}e^{-\mathcal{H}(M)})$, where $M$ is the set of configurations that disagree with your ground configuration at only finitely many points, and have energy below your chosen threshold. Such a solution identifies the points where the two modified pressure of the two ground states are equal, and so both have corresponding Gibbs states: the coexistence line. This equation is difficult to solve explicitly, so we instead solve up to the leading term in the expansion in $e^{-\varepsilon_{i}}$, where $\varepsilon_{i}\in\mathcal{E}$ and $\mathcal{E}$ is the set of possible excitation energies. Notate equation up to leading term by $\approx$. We find the coefficients using the cluster expansion. Because we only consider the leading order term, which consist of single polymers, the coefficient in $\dot{P}^{G}$ simplifies to the density of minimal excitations. In the following calculations, we will use $\beta$ for $1/T$, to simplify the notation.
The $E$ configuration only has one kind of minimal excitation, flipping a single hexagon. This changes 6 vertices in state $E$ to 6 vertices in state $C$, changing the energy by $1-f$. The perturbation of the Hamiltonian assigns $E$ 0 energy, so that term disappears.
The $H$ configuration has one of two possible minimal excitations, depending on the choice of parameters. Flipping one of the empty hexagons which form 1/3 of the configuration replaces 6 $H$ vertices with $F$ vertices, changing the energy by $6(f-h)$. Flipping a full hexagon replaces 6 $H$ vertices with $C$ vertices, changing the energy by $6(1-f-h)$. Combining this information, we solve
$e^{6(1-f)\beta}\approx\frac{1}{3}e^{6f\beta-6h}+\frac{2}{3}e^{6(1-f)\beta-6h}+h$
@@
$h\approx\frac{1}{3}e^{6(1-f)\beta}-\frac{1}{3}e^{6f\beta}$
Therefore $h>0$ when $f<1/2$ and $h>0$ when $f>1/2$.
At $f=1/2$, there is a symmetry between the $E$ and $H$ excitation expansions. For any minimal excitation, flipping the value of hexagons in two sublattices replaces vertices in state $E$ with $H$ and vice versa, while vertices in state $F$ are replaced with $C$, again vice versa. Therefore if $e_{H}=e_{E}$ and $e_{f}=e_{C}$, the multiplicities and energies of excitations around $E$ and $H$ will be the same, the pressures will be the same, and so by symmetry the line of the phase transition must pass through $f=1/2$ at $h=0$.
The behavior along $F-C$ curve is the same, by inversion symmetry.
Theorem 3.
There exists $T_{0}$ such that $\forall T<T_{0}$, along the $E-F$ transition line, if $e_{H}>e_{C}$, the $E$ state dominates and if $e_{H}<e_{C}$ the $F$ state dominates.
Proof.
[…]
$e^{(1-h)\beta}\approx e^{h\beta-f}+f$
@@
$f\approx e^{(1-h)\beta}-e^{h\beta}$
so $f>0$ when $h<1/2$ and $f>0$ when $h>1/2$. By the symmetry of the two ground states, we know this curve must be odd around the equality point, and thus passes exactly through 0 there. ∎
@@ 3.3 Non-Peierls Lines
Now we handle two of the three non-Peierls infinitely degenerate transition lines. We can’t use Pirogov–Sinai here, so we will instead use disagreement percolation to prove uniqueness, then apply reflection positivity to understand the behavior. Our argument relies on the following uniqueness condition [BM94]:
Theorem (Disagreement Percolation).
Fix a particular Markov field, and notate $N_{i}$ as the set of neighbors of a single site, $d(\cdot,\cdot)$ as the variational distance, and $Y_{i}(\cdot,\eta)$ as the one-site distribution given its neighbors. If $p_{i}<p_{c}$, where $p_{i}=\max_{\eta,\eta^{\prime}\in\{0,1\}^{N_{i}}}d(Y_{i}(\cdot,\eta),Y_{i}(\cdot,\eta^{\prime})$ and $p_{c}$ is the critical site percolation threshold for the given lattice, then there is exactly one Gibbs measure induced by the Markov field.
Theorem 4.
Along the $E-C$ transition line, if $e_{F}\geq e_{H}$, there is a unique Gibbs state at all temperatures. By inversion symmetry, the same is true along the $F-H$ line when $e_{C}\geq e_{E}$.
@@
Theorem 5.
Along the $E-C$ transition line, if $e_{F}\geq e_{H}$ then no configuration dominates at any temperature.
Proof.
If a Gibbs state is unique, it must also be translation-invariant: any translation of a Gibbs state is still a Gibbs state, so if there’s exactly 1, it must be equal to all it’s translations. Therefore, any dominating configuration must be invariant for any translation preserving the lattice. However, the only translation-invariant configurations are $E$ or $F$. The $F$ configuration has maximal energy, and so is minimally likely among all configurations. If we assume the $E$ configuration dominates, then the probability of any face being in state $E$ must go to unity as $T\to 0$. However, any face whose neighbors are all in state $E$ can have its value flipped without increasing the energy, so that configuration must be equally likely as the $E$ configuration, conditional on empty neighbors. Thus, the probability of a filled hexagon cannot go to 0: it must be no less than $1/2$ the probability that all its neighbors are empty. If the probability that all its neighbors are empty goes to 0, the probability of a filled neighbor cannot go to 0, and we have a contradiction by translation invariance.
Although we lack domination, at low temperatures the set of allowed configurations are still constrained with high probability. However, instead of approximating a particular configuration, they are constrained to avoid particular vertex states. Because the Gibbs measure is unique and the space of measures is sequentially compact, the unique infinite Gibbs measure is achieved by any sequence of local measures with infinite limiting radius. Thus, we can restrict our attention to measures defined on the rectangular torus. A reflection is defined by a partition of the torus into two halves, by a pair of horizontal or vertical lines, such that reflecting across the lines preserves the lattice. We will use the “closed half-space” to refer to the set of faces intersecting of one of these two halves. The particular reflections we’re considering here are represented in Figure 8.
@@
The first equation follows immediately from $\theta$ invariance of the Hamiltonian under both kinds of reflections. For horizontal reflections through the faces, the locality of the model immediately implies the second relation holds; see the argument in [HP22]. For vertical reflections through sites and bonds, we use the following statement: if the Hamiltonian can be realized as $A+\theta(A)+\sum_{i}C_{i}\theta(C_{i})$, with $A,C_{i}\in A^{+}$, the induced measure is reflection positive [FV17]. Each vertex internal to one of the half-spaces has an associated potential $A\in A^{+}$, and a corresponding vertex in the other half with the potential in $\theta(A)$. However, the vertices along the border do not have potentials in $A^{+}$, so representing the associated potential in these terms is nontrivial. To achieve this, we parametrize all functions $C_{a,b,c,d}\in A^{+}$ that depend only on a pair of boundary hexagons by the energy they assign to each of the four possible configurations. Then, we combine them to look at the associated functions on triplets in $A^{+}\theta(A^{+})$ and $A^{+}+\theta(A^{+})$. These functions are represented visually in Figure 8.
Figure 7: Local energy functions considering two boundary hexagons, and the induced functions on three boundary hexagons, two in each half space, with the top hexagon in both closed half-spaces. With these we construct indicator functions for each vertex state
Figure 8: Horizontal reflections (blue) and vertical reflections (red) define blocks containing one full vertex

### v2

@@ 3 Proofs of the the Main Theorems
Before we make any specific arguments, first we should note the symmetry of the model. Flipping all the spins interchanges $E$ vertices for $F$ vertices, $C$ vertices for $H$ vertices, and vice versa. Thus, flipping the spins while swapping the energies assigned to the respective vertex states maintains the assignment of probabilities. With this in mind, any statement that concerns particular vertex states applies just as well to its “twin” under spin-flip symmetry. We will frequently make an argument for one state, then apply the same to its twin by spin-flip symmetry.
@@
Using techniques from Pirogov–Sinai, we address the large regions dominated by a particular vertex state first, in subsection 3.1. Next, with Slawny’s technique we handle the position of the coexistence curves between these regions in subsection 3.2. Finally, with disagreement percolation and reflection positivity we determine the behavior along the degeneracy lines for which Pirogov–Sinai techniques are inapplicable, in subsection 3.3.
@@ 3.1 Applying Pirogov–Sinai
We begin by describing the ground states of the model, which define the regions in Fig. 4. Given two vertex states $A,B$, we will refer to the points where $e_{A}=e_{B}$ and both achieve the minimal vertex energy as the ABA-B degeneracy line.
Lemma 1.
Along the HFH-F, ECE-C, or HCH-C degeneracy lines, there are infinitely many ground configurations. At all other points, there are finitely many ground configurations, all of which are periodic.
@@
First, we consider the case where exactly one vertex state has minimal energy. If a configuration with only that vertex state is possible, it will be minimal. For the $F$ and $E$ vertex states, minimal energy is achieved by the full and empty configurations respectively; any other configuration must have other vertex states, and so the ground configuration is unique and translation invariant. For the $H$ and $C$ vertex states, we start with a single vertex. This could have three distinct assignments, as the $C$ state is realizable by three different three-hexagon configurations. However, given a trio of three hexagons and state $C$, there is only one configuration on its neighbors such that every vertex is in state $C$. By induction, this uniquely determines a periodic ground configuration given one of our original choices, so there are three periodic ground configurations. Along the EHE-H, EFE-F, and CFC-F degeneracy lines, the minimal vertex states cannot coexist, i.e. neighboring vertices cannot have distinct minimal vertex states, they must be separated by a third vertex that is in a non-minimal state, as demonstrated in Fig. 6. Therefore, all minimal configurations must have only one vertex state, so the set of minimal configurations is the union of the set on the two neighboring regions, and as such is finite and periodic.
Figure 6: Each of the three pairs of vertex states, EH,EF,CFE-H,{E-F},C-F which cannot be directly adjacent, but must have a third vertex state between them
Along the HFH-F, ECE-C, and HCH-C boundary lines, minimal vertex states can coexist. Thus there are infinitely many energy-preserving modifications to particular ground configurations (for example, given the empty configuration, along the ECE-C line, one can fill in any single hexagon), so the set of ground configurations is infinite. These ground configurations are also not in general periodic: the lower half-plane may have one vertex state, while the upper may have the other.
@@
Lemma 2.
All Hadwiger models satisfy the Peierls condition, except on the HFH-F, ECE-C, or HCH-C degeneracy lines.
@@
An $m$-potential is a function on $\omega\in\{-1,1\}^{S}$ that can be written as $\sum_{U\subset S}f_{U}(\omega)$, where each $f_{U}$, called a “potential”, depends only on elements in $U$, and there exists a configuration $\omega$ such that $f_{U}$ achieves its minimal value for all $U$. As described above, any Hadwiger Hamiltonian can be represented as the sum over the vertices of terms that depend only on the neighboring hexagons, and may have every vertex in a minimum-energy state, so the Hadwiger models are $m$-potentials. If an $m$-potential’s set of minimal configurations is finite, it satisfies the Peierls condition [FV17]. As just demonstrated, outside the HFH-F, ECE-C, or HCH-C degeneracy lines the set of minimal configurations is finite, so at those points the Peierls conditions is satisfied.
Along the HFH-F, ECE-C, and HCH-C degeneracy lines, there are infinitely many ground configurations; these are achieved by starting from one ground configuration and flipping any of the infinitely many hexagons that do not increase the energy. Thus, they do not satisfy the Peierls condition and have nonzero entropy at 0 temperature. In particular, the ground states of HFH-F and ECE-C are the allowed states of the hard hexagon model, while the ground states of the HCH-C model are ground states of the frustrated antiferromagnetic Ising model, or equivalently, random lozenge tilings [G21]. ∎
@@
Proof of Theorem 3.
@@
Zahradnik [Z84] states for any periodic Peierls model with finite ground configurations $\exists T_{0}$ such that $\forall T<T_{0}$, every periodic Gibbs state is a linear combination of extremal Gibbs states dominated by a periodic ground configuration. Dobrushin and Shlosman [DS85] state that for 2-dimensional models, under the same conditions, every Gibbs state is periodic. Therefore, for any 2-dimensional periodic Peierls model with finite ground configurations $\exists T_{0}$ such that $\forall T<T_{0}$, every Gibbs state is a linear combination of those dominated by the ground states. The model always satisfies the other conditions, so at every Peierls point at sufficiently low temperatures, the space of Gibbs state is exactly those dominated by the ground states.
3.2 Coexistence Curves
@@
At fixed low temperature, there are three coexistence curves to consider, the latter two of which are equivalent under spin-flip symmetry: EFE-F, EHE-H, and CFC-F. Zahradnik tells us that coexistence curves approach the zero-temperature degeneracy lines, but does not tell us from which direction they approach it, or whether they achieve it at some nonzero temperature. There is a technique by Slawny [S87] to calculate the asymptotic low-temperature positions of these coexistence curves explicitly, by calculating at each point which of the ground configurations has the lowest-energy or highest-multiplicity minimal excitation. We use this technique in a slightly unconventional way. Normally Slawny’s technique is used to determine the location of a low-temperature triple point, where three coexistence curves meet, relative to a Peierls zero-temperature triple point. No Peierls triple point exists here, but we can still determine the relative location of the coexistence curve by calculating 1-dimensional “slices” perpendicular to each point along the zero-temperature degeneracy line. In each slice, there is a nonzero $T$ such that the transient terms are smaller than the asymptotic term in the expansion we describe below. From Slawny we know the terms vary continuously, so on a closed region there will be a nonzero minimum, such that the overall curve (which by implicit function theorem must be continuous) approximates the asymptotic curve.
Proof of Theorem 4.
@@
Along this line, the $E$ and $H$ vertex states have energy 0, and moving along the line changes the energy of $F$ and $C$. The space of models lies on a sphere, but we’re interested only in the asymptotic shape and position relative to the degeneracy lines, so we can change coordinates at will as long as it does not affect relative position. Project to a plane to simplify coordinates, so along the degeneracy line $e_{F}=f\in[\epsilon,1-\epsilon]$ and $e_{C}=1-f$. Moving perpendicular to this line, $e_{H}=h$. Following Slawny, we can treat the three symmetric ground configurations as a single ground configuration for the purposes of the calculation.
@@
$\dot{P}^{G}(\beta\mathcal{H}_{0}+\mathcal{H})-e_{G}(\mathcal{H}^{\prime})=\dot{P}^{G^{\prime}}(\beta\mathcal{H}_{0}+\mathcal{H}^{\prime})-e_{G^{\prime}}(\mathcal{H}^{\prime})$ (4)
where $\dot{P}^{G}$ is the pressure of the gas of excitations from the ground configuration $G$ with respect to a given Hamiltonian, $\mathcal{H}_{0}$ is the Hamiltonian you’re expanding around, $\mathcal{H}^{\prime}$ is the perturbation to the Hamiltonian, and $e_{G}$ is the energy density of $G$ with respect to a given Hamiltonian, equivalent to the energy the Hamiltonian assigns to the associated vertex state. By “pressure of the gas of excitations”, we mean $\log(\sum_{m\in M}e^{-\mathcal{H}(M)})$, where $M$ is the set of configurations that disagree with your ground configuration at only finitely many points, and have energy below your chosen threshold. Such a solution identifies the points where the modified pressures of the two ground states are equal, and so both have corresponding Gibbs states: the coexistence curve. This equation is difficult to solve explicitly, so we instead solve up to the leading term in the expansion in $e^{-\varepsilon_{i}}$, where $\varepsilon_{i}\in\mathcal{E}$, the set of possible excitation energies. Denote equation up to leading term by $\approx$. We find the coefficients using the cluster expansion. Because we only consider the leading order term, which consist of single polymers, the coefficient in $\dot{P}^{G}$ simplifies to the density of minimal excitations. In the following calculations, we will use $\beta$ for $1/T$, to simplify the notation.
The $E$ configuration only has one kind of minimal excitation, flipping a single hexagon. This changes 6 vertices in state $E$ to 6 vertices in state $C$, changing the energy by 1f1-f. The perturbation of the Hamiltonian assigns $E$ 0 energy, so that term disappears.
The $H$ configuration has one of two possible minimal excitations, depending on the choice of parameters. Flipping one of the empty hexagons which form 1/3 of the configuration replaces 6 $H$ vertices with $F$ vertices, changing the energy by 6(fh)6(f-h). Flipping a full hexagon replaces 6 $H$ vertices with $C$ vertices, changing the energy by 6(1fh)6(1-f-h). Combining this information, we solve
$e^{6(1-f)\beta}\approx\frac{1}{3}e^{6f\beta-6h}+\frac{2}{3}e^{6(1-f)\beta-6h}+h.$ (5)
@@
$h\approx\frac{1}{3}e^{6(1-f)\beta}-\frac{1}{3}e^{6f\beta}.$ (6)
Therefore $h>0$ when $f<1/2$ and $h<0$ when $f>1/2$.
At $f=1/2$, there is a symmetry between the $E$ and $H$ excitation expansions. For any minimal excitation, flipping the value of hexagons in two sublattices replaces vertices in state $E$ with $H$ and vice versa, while vertices in state $F$ are replaced with $C$, again vice versa. Therefore if $e_{H}=e_{E}$ and $e_{F}=e_{C}$, the multiplicities and energies of excitations around $E$ and $H$ will be the same, the pressures will be the same, and so by symmetry the curve of the phase transition must pass through $f=1/2$ at $h=0$. This implies the coexistence curve passes through the point at $f=1/2$, with the $E$ phase extending over the line when $f>1/2$ and the $H$ phase extended over the line when $f<1/2$.
The behavior along CFC-F curve is the same, by spin-flip symmetry.
Proof of Theorem 5.
Again we project to a plane, $F$ and $E$ both have energy 0, $e_{h}=h\in[-1+\epsilon,1-\epsilon]$, and $e_{C}=1-h$ in $1+\epsilon,1-\epsilon]$. Moving perpendicular to this line, $e_{F}=f$. In the $E$ configuration, flipping an empty hexagon replaces 6 $E$ vertices with $C$ vertices, changing the energy by 1h1-h, while in the $F$ configuration, flipping a full hexagon replaces 6 $F$ vertices with $H$ vertices, changing the energy by hfh-f. Thus, we solve
$e^{(1-h)\beta}\approx e^{h\beta-f}+f$ (7)
@@
$f\approx e^{(1-h)\beta}-e^{h\beta}$ (8)
so $f>0$ when $h<1/2$ and $f<0$ when $h>1/2$. By the symmetry of the two ground states, we know this curve must be odd around the equality point, and thus passes exactly through 0 there. ∎
@@ 3.3 Non-Peierls Lines
Now we handle two of the three non-Peierls infinitely degenerate degeneracy lines. We can’t use Pirogov–Sinai here, so we will instead use disagreement percolation to prove uniqueness, then apply reflection positivity to understand the behavior. Our argument relies on the following uniqueness condition [BM94]:
Theorem 8 (Disagreement Percolation).
Fix a particular Markov field, and denote $N_{i}$ as the set of neighbors of a single site, $d(\cdot,\cdot)$ as the variational distance, and $Y_{i}(\cdot,\eta)$ as the one-site distribution given its neighbors. If $p_{i}<p_{c}$, where $p_{i}=\max_{\eta,\eta^{\prime}\in\{-1,1\}^{N_{i}}}d(Y_{i}(\cdot,\eta),Y_{i}(\cdot,\eta^{\prime})$ and $p_{c}$ is the critical site percolation threshold for the given lattice, then there is exactly one Gibbs measure induced by the Markov field.
Corollary 9.
Along the ECE-C degeneracy line, if $e_{F}\geq e_{H}$, there is a unique Gibbs state at all temperatures. By spin-flip symmetry, the same is true along the FHF-H line where $e_{E}\geq e_{C}$.
@@
Proof of Theorem 6.
From Corollary 9 above, we know that the Gibbs state is unique. If a Gibbs state is unique, it must also be translation-invariant. Therefore, any dominating configuration must also be translation invariant. However, the only translation-invariant configurations are $E$ and $F$. The $F$ configuration has maximal energy and so is minimally likely among all configurations. If we assume the $E$ configuration dominates then the probability of any face being in state $E$ must go to unity as $T\to 0$. However, any face whose neighbors are all in state $E$ can have its value flipped without increasing the energy, so that configuration must be equally likely as the $E$ configuration, conditional on empty neighbors. Thus, the probability of a filled hexagon cannot go to 0: it must be no less than $1/2$ the probability that all its neighbors are empty. If the probability that all its neighbors are empty goes to 0, the probability of a filled neighbor cannot go to 0, and we have a contradiction by translation invariance. ∎
Although we lack domination, at low temperatures the set of allowed configurations is still constrained with high probability. However, instead of approximating a particular configuration, they are constrained to avoid particular vertex states. Because the Gibbs measure is unique and the space of measures is sequentially compact, the unique infinite Gibbs measure is achieved by any sequence of local measures with infinite limiting radius. Thus, we can restrict our attention to measures defined on the rectangular torus. A reflection is defined by a partition of the torus into two halves, by a pair of horizontal or vertical lines, such that reflecting across the lines preserves the lattice. We will use the “closed half-space” to refer to the set of faces intersecting of one of these two halves. The particular reflections we’re considering here are represented in Fig. 9.
Figure 8: (left)Local energy functions considering two boundary hexagons, and the induced functions on three boundary hexagons, two in each half space, with the top hexagon in both closed half-spaces. With these we construct indicator functions for each vertex state
Figure 9: Horizontal and vertical reflections define blocks containing one full vertex
@@
The first equation follows immediately from $\theta$ invariance of the Hamiltonian under both kinds of reflections. For horizontal reflections through the faces, the locality of the model immediately implies the second relation holds; see the argument in [HP22]. For vertical reflections through sites and bonds, we use the following statement: if the Hamiltonian can be realized as $A+\theta(A)+\sum_{i}C_{i}\theta(C_{i})$, with $A,C_{i}\in A^{+}$, the induced measure is reflection positive [FV17]. Each vertex internal to one of the half-spaces has an associated potential $A\in A^{+}$, and a corresponding vertex in the other half with the potential in $\theta(A)$. However, the vertices along the border do not have potentials in $A^{+}$, so representing the associated potential in these terms is nontrivial. To achieve this, we parametrize all functions $C_{a,b,c,d}\in A^{+}$ that depend only on a pair of boundary hexagons by the energy they assign to each of the four possible configurations. Then, we combine them to look at the associated functions on triplets in $A^{+}\theta(A^{+})$ and $A^{+}+\theta(A^{+})$. These functions are represented visually in Fig. 9.